OpenStudy (anonymous):

Jolene is putting trim around the edge of a triangle head scarf. The scarf from a right triangle with legs that measures 15 inches each. Find how much trim Jolene needs to the nearest tenth of an inch

OpenStudy (anonymous):



OpenStudy (anonymous):

Since it's a right triangle, you can use pythagorean theorem to find the hypotenuse. a^2 + b^2 = c^2

OpenStudy (anonymous):

So it would be 15^2 + 15^2 = c^2 225+225=c^2 /450=c 21.21=c Right?

OpenStudy (anonymous):

Exactly! Since you have now identified all three sides, you can now find its perimeter!

OpenStudy (anonymous):

Thank you. :)